A+T Williamson Company is making adjusting entries for the year ended December 31 of the current year. In developing information for the adjusting entries, the accountant learned the following:

  1. A two-year insurance premium of $3,120 was paid on October 1 of the current year for coverage beginning on that date. The bookkeeper debited the full amount to Prepaid Insurance on October 1.
  2. At December 31 of the current year, the following data relating to Shipping Supplies were obtained from the records and supporting documents.

   Shipping supplies on hand, January 1 of the current year $17,500

Purchases of shipping supplies during the current year 62,500

Shipping supplies on hand, counted on December 31 of the current year 13,000

3. What amount should be reported on the current year's income statement for Insurance Expense? For Shipping Supplies Expense?

What amount should be reported on the current year's balance sheet for Prepaid Insurance? For Shipping Supplies?
Answer #1

3) Income statement

Insurance expense (3120/24*3) = 390

Shipping supplies expense (17500+62500-13000) = 67000

4) Balance Sheet

Prepaid insurance = 3120/24*21 = 2730

Shipping supplies = 13000